1. What is the percent ionization of a 1.38 mol/L weak base if its Kb = 2.7 x 10-6?
  2. What is Kb for N2H4 if the pH of a 0.158M solution of N2H5Cl is 4.5?
1
Expert's answer
2020-10-27T12:41:19-0400

1. Kb=a^2xC

a=✓(Kb/C)=✓(2.7x10-6/1.38)= 0,00139875721. = 0.1398 %

2. [H+]=antipH=3.16227766017E-05

a=3.16227766017E-05/0.158 = 0,000200144156

Kb=a^2xC = 0,000200144156^2 x 0.158 = 6,32911395e-9.